# Dynamical scheme for computing the mass parameter of a system in a medium

### v1

@@ I Introduction
Effective models are indispensable tools in physics that inevitably emerge for complex systems. With their help, one can connect a physical phenomenon taking place across various scales. Whenever we change degrees of freedom and construct a new effective model operating at larger length scales, input from the underlying model is required. In this way, we obtain a hierarchy of models spanning from macroscopic approaches through mesoscopic ones up to microscopic ones. As an example, relevant to this paper, let us consider the problem of constructing an effective model of a rotating neutron star. At the largest scales, it can be modeled by means of hydrodynamical approaches Graber et al. (2017); Shibata et al. (2017). To explain observed phenomena of neutron star glitches (sudden decreases of rotation period), one needs to revert to a multifluid picture, with the extra assumption that one of the fluids is in a superfluid state Haskell and Melatos (2015); Khomenko and Haskell (2018). Besides the equation of state, such models require various macroscopic parameters to be provided, like viscosities, conductivities, and mutual friction parameters between the fluid components. Going down in the hierarchy, by increasing the resolution, we enter the regime of mesoscopic models, where we start to resolve quantized vortices present in the superfluid component, which coexists with the so-called impurities created by the other fluid component Howitt et al. (2020); Cheunchitra et al. (2024). Precisely, neutron matter forms the superfluid component, filled out with vortices, while protons form various structures, which for the majority of the inner crust have spherical shapes. Electrons, which are essential for charge neutrality, are well approximated by their uniform distribution. At this level, one usually refers to the Vortex Filament Model (VFM), which was originally developed for modeling vortices in superfluid helium Tsubota et al. (2013). It is a quasiclassical model, where one writes Newton’s equations for quantum vortices (presented as filaments in 3D or points in 2D) that interact with all other elements of the system. The drag force exerted on vortices by the impurities is the origin of the mutual force that appears in the hydrodynamical model. Increasing further the resolution, we start to resolve the building blocks of the neutron’s star matter: neutrons, protons, and electrons (and possibly other elementary particles deep in the star’s core). At this level, one needs to revert to the microscopic description grounded in many-body quantum mechanics. The quantized vortices turn out to be a collective flow of neutrons around some line in space, while impurities correspond to regions where protons (together with neutrons) concentrate into various structures Negele and Vautherin (1973); J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018), usually into nuclei-like objects, but more complex structures are also allowed Caplan and Horowitz (2017). The effective interaction between the flow of neutrons and the bound state of protons is the origin of the drag force that enters the mesoscopic models.
@@
where $Q$ is the selected collective variable. Such a form is achievable once the adiabatic limit can be recovered in the system, which means that the terms describing couplings to other degrees of freedom can be neglected Baranger and Vénéroni (1978). The first term is the kinetic term that is quantified by the mass parameter $M_{Q}$, and $V$ stands for the potential energy. The choice of the collective variable is motivated by physical intuition, although a more precise definition is also possible. In the latter case, one searches for the collective path (or space), which is maximally decoupled from the other degrees of freedom (see Washiyama et al. (2024); Wen and Nakatsukasa (2022); Nakatsukasa et al. (2016) and references therein). This can be important in the case where the choice of collective variables is not obvious, as in the case of exotically deformed nuclei forming the pasta phase. Whatever method is applied to identify $Q$, it implicitly assumes that the adiabatic limit can be achieved, in which the motion of the system is governed by Hamiltonian (1).
In this paper, we present a general-purpose method for extracting mass parameters for nuclear systems by means of the time-dependent density functional theory approach (TDDFT). While there are methods for extracting the effective mass parameters, most of them rely on various assumptions, and they are typically designed to be applicable for extracting the mass parameters in a vacuum (see, e.g., the review paper Li and Vretenar (2022)). The presented approach is free of these limitations, and we will demonstrate it for the nuclear impurity immersed in superfluid neutron liquid, i.e., for the conditions relevant to the neutron star crust. The calculations will be delivered for mass parameters associated with translational motion and quadrupole vibrations. The selected collective variables are schematically presented in Fig. 1.
@@ II Method
To make the problem more concrete, let us consider a Wigner crystal of neutron-rich nuclei in the crust of a neutron star. These nuclei can oscillate within the crystal, as illustrated in Fig. 1(a). The collective variable $Q$ can then be associated with the impurity’s position $\bm{R}_{\textrm{CM}}$, which, in a microscopic description, corresponds to the center of mass of the protons. The potential energy in this system is primarily due to Coulomb interactions between impurities. Even in this relatively simple setting, determining the effective mass of an impurity immersed in neutron matter remains a challenging task. A variety of approaches have been developed to address this problem Magierski (2004); Magierski and Bulgac (2004); Martin and Urban (2016); N. Chamel, D. Page and S. Reddy (2013); Pęcak et al. (2024). The Wigner crystal is also expected to be threaded by quantum vortices. Work G. Wlazłowski, K. Sekizawa, P. Magierski, A. Bulgac and M. McNeil Forbes (2016) demonstrated that the interaction between an impurity and a vortex can induce a quadrupole deformation of the impurity, Fig. 1(b). This suggests that the quadrupole moment $Q_{20}$ could also serve as a relevant collective variable in the effective model. It is also possible that higher-order nuclear multipole moments play a significant role.
The shape of the impurity can be parametrized using an expansion in spherical harmonics $Y_{\lambda\mu}(\theta,\phi)$
$R(\theta,\phi)=R_{0}\left(1+\sum_{\lambda=0}^{\infty}\sum_{\mu=-\lambda}^{\lambda}\alpha_{\lambda\mu}Y_{\lambda\mu}(\theta,\phi)\right).$ (2)
where $R_{0}$ is the radius of the undeformed impurity, and each deformation parameter $\alpha_{\lambda\mu}$ is associated with a corresponding mass parameter. In this expansion, $\lambda=1$ corresponds to the dipole mode, which relates to the impurity’s position, while $\lambda=2$ corresponds to the quadrupole mode. These two modes have been identified as relevant collective variables for the mesoscopic model and are therefore the focus of investigation in this work.
@@
It is worth mentioning that there are other methods to extract collective mass parameters that are typically defined in a vacuum Li and Vretenar (2022). As a result, no existing method works for nuclei in a medium, such as the inner crust of a neutron star. One example is the adiabatic approximation to the time-dependent Hartree-Fock-Bogoliubov approach (ATDHFB), commonly used to analyze large amplitude collective motion Baran et al. (2011). The fundamental idea behind this method is that the collective motion of the whole system is slow compared to the independent movement of each of the nucleons that the system consists of. This approach has various modifications, including the popular cranking approximation Girod and Grammaticos (1979). All these approaches extract the mass parameter from quasiparticle wave function, corresponding to static solutions of time-independent HFB equations ($i\hbar\partial/\partial t\rightarrow E_{\eta,q}$ in Eq. (9)). Such methods encounter difficulties in cases of more complex problems, like nuclei immersed in the neutron matter. In such a case, one cannot easily split the quasiparticle wave functions into two sets, where one represents nucleons belonging to nuclei and the other represents nucleons belonging to the background. The approach presented here is free of these difficulties, as it relies only on the observables constructed from densities. Therefore, the method is universal and can be applied to the region of the inner crust of a neutron star to extract the effective mass parameters associated with various collective degrees of motion.
@@
The method described above is generic and does not depend on specific functional details. In this study, we use the Brussels-Montreal class of Skyrme-type functionals, which are particularly well suited for studies in the context of neutron stars. These functionals are developed using not only experimental data biased on finite-size nuclei but also many-body results Cao et al. (2006) that account for infinite nuclear matter. In this way, the bulk properties are incorporated into the model. The Brussels-Montreal-Skyrme functional has a generic form:
[…]
@@
$h_{q\sigma\sigma}=-\bm{\nabla}\frac{\delta E}{\delta\tau_{q}}\bm{\nabla}+\frac{\delta E}{\delta\rho_{q}}-\frac{i}{2}\left\{\frac{\delta E}{\delta\bm{j}_{q}},\bm{\nabla}\right\},\$ (17)
[…]
@@ III.1 Center of mass oscillations
[…]
$\begin{split}U_{\textrm{p}}=\frac{\delta E}{\delta\rho_{p}}=U_{\textrm{p},\textrm{BSk}}+k\left(Z_{\textrm{CM}}(t)-Z_{0}\right)\frac{z}{N_{p}}.\end{split}$ (22)
@@
In order to examine the system described above, we did systematic studies for average nuclear densities $\bar{\rho}$ ranging from 0.002 $0.002\textrm{~{}fm}^{-3}$ to 0.051 $0.051\textrm{~{}fm}^{-3}$, where the average density is defined as $\bar{\rho}=(N_{n}+N_{p})/V_{WS}$, ie., number of nucleons (protons and neutrons) per Wigner-Seitz cell volume J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018). We have generated initial states, as shown in Fig. 1, by means of solving the static variant of problem Eq. (9). The problem was solved on a spatial mesh of size $N_{x}\times N_{y}\times N_{z}=32\times 32\times 64$ with lattice spacing $\Delta x=\Delta y=\Delta z=1.25$ fm. The associated cutoff energy is fixed through the value of resolution to $E_{\mathrm{cut}}=\hbar^{2}\pi^{2}/(2m_{n}\Delta x^{2})\approx 130$ MeV. In the computation, we have used W-BSK Toolkit, described extensively in Ref. Pęcak et al. (2024), and available as an open-source package via webpage WBS . As a side remark, we note that the toolkit neglects contributions arising from the spin-orbit part of the nuclear interaction, which in turn allows the execution of simulations in sufficiently large volumes relevant to nuclear problems in the context of neutron star studies. Next, we have performed time-dependent simulations with modified mean-field potential Eq. (22) with the initial displacement $Z_{0}=1$fm.
@@
Figure 2: a) Position of the center of mass of protons as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8).
@@
We applied the fitting procedure to extract the effective parameters of the impurity. Fig. 2 shows the center of mass oscillations for different values of control parameter $k$, for selected nuclear density $\bar{\rho}=0.015\textrm{~{}fm}^{-3}$. As expected, the extracted frequencies of oscillations follow Eq. (8). One can clearly see that the intercept in this case is $\omega_{Z_{\textrm{CM}}}=0$. This feature of the center of mass collective degree of freedom is also expected. Since the Hamiltonian is spatially symmetric, i.e., it is invariant under the operation $\bm{r}\rightarrow\bm{r}+\bm{r}_{0}$, shifting the nucleus does not change the energy of the system. Nevertheless, the translational symmetry is broken: the nuclear cluster can form at any place, which is set randomly. The $\omega_{Z_{\textrm{CM}}}=0$ mode is then analogous to a zero-energy Goldstone mode associated with the breaking of translational symmetry. This mode does not represent oscillation but rather a shift with a constant velocity that moves the ground state across all degenerate quantum states that are also solutions of the Hamiltonian.
@@
Figure 3: The effective mass $M_{\textrm{CM}}$ (in units of neutron mass $m_{n}$) as a function of mean nuclear density $\bar{\rho}$ extracted for small amplitude $A=1$ fm (orangle triangles) and large amplitude $A=1$ fm (green squares) oscillations. We compare the results of the proposed method (orange triangles) with the effective mass obtained recently using a different dynamical scheme (blue diamonds) Pęcak et al. (2024).
@@
Figure 4: a) The damped oscillations of the center of mass of protons for selected density $\bar{\rho}=0.019\textrm{~{}fm}^{-3}$ (we show that point in panel c), and b) associated absolute value of velocity as a function of time. The velocity is compared to Landau’s velocity. In panel c), we show the extracted damping coefficient $\gamma_{\textrm{CM}}$ as a function of average nuclear density.
@@
where $\Delta_{n}$ and ${k_{\mathrm{F}}}_{n}$ are the bulk values of the pairing field and the Fermi wave-vector of neutrons. Note that this quantity is defined for a uniform system, and it is not obvious how well it works for such a non-uniform system as we study. For increased amplitude, we observed damped oscillations as shown in Fig. 4a). From fitting to the Eq. (19) we can accurately reproduce the behaviour of damped oscillations (orange solid line) and extract the damping parameter $\gamma_{\textrm{CM}}$. We investigate the absolute value of the velocity of the impurity in Fig. 4b). One can clearly connect the damping with exceeding the Landau velocity by the impurity. The damped oscillator model starts to underestimate the oscillations’ amplitude once the maximum velocity drops significantly below the Landau velocity. A similar mechanism was observed in recent work Pęcak et al. (2024) where the breaking of Cooper pairs was one of the mechanisms of energy dissipation. The last panel of Fig. 4c) shows the damping parameter $\gamma_{\textrm{CM}}$ as a function of average nuclear density in the inner crust. This is a monotonically increasing function of the density. In the limit of low density $\bar{\rho}\rightarrow 0$, it corresponds to a nucleus oscillating in the vacuum. Obviously, in this regime, there are no degrees of freedom where the energy can be dissipated, thus no damping is present. Once the medium gets denser, we see that the damping coefficient increases as well. This effect comes from the investigation of Eq. (23) where the pairing field $\Delta_{n}$ vanishes in high densities Cao et al. (2006), while the wave vector in the denominator $k_{\mathrm{F}}$ grows. Note that in this work, we consider only the ${}^{1}S_{0}$ pairing channel, and neglect other channels for the pairing, such as higher angular momentum channels ${}^{3}P_{1}$, which is believed to dominate in higher densities.
The case of oscillations with amplitude $A=10$ fm cannot be considered small, as the amplitude becomes comparable to, or even exceeds, the size of the protonic cluster itself, whose radius ranges from $5$ to $10$ fm depending on the density. Nevertheless, extracting the mass parameter from these data still yields reasonable results, see square markers in Fig. 3. Although the method consistently overestimates $M_{\textrm{CM}}$ by approximately $10-15$%, the qualitative behavior is well captured. This discrepancy can be viewed as an estimate of the upper bound on the method’s uncertainty.

### v2

@@ I Introduction
Effective models are indispensable tools in physics that inevitably emerge for complex systems. With their help, one can connect a physical phenomenon taking place across various scales. Whenever we change degrees of freedom and construct a new effective model operating at larger length scales, input from the underlying model is required. In this way, we obtain a hierarchy of models spanning from macroscopic approaches through mesoscopic ones up to microscopic ones. As an example, relevant to this paper, let us consider the problem of constructing an effective model of a rotating neutron star. At the largest scales, it can be modeled by means of hydrodynamical approaches Graber et al. (2017); Shibata et al. (2017). To explain observed phenomena of neutron star glitches (sudden decreases of rotation period), one needs to revert to a multifluid picture, with the extra assumption that one of the fluids is in a superfluid state Haskell and Melatos (2015); Khomenko and Haskell (2018). Besides the equation of state, such models require various macroscopic parameters to be provided, like viscosities, conductivities, and mutual friction parameters between the fluid components. Going down in the hierarchy, by increasing the resolution, we enter the regime of mesoscopic models, where we start to resolve quantized vortices present in the superfluid component, which coexists with the so-called impurities created by the other fluid component Howitt et al. (2020); Cheunchitra et al. (2024). Precisely, neutron matter forms the superfluid component, filled out with vortices, while protons form various structures, which for the majority of the inner crust have spherical shapes. Electrons, which are essential for charge neutrality, are well approximated by their uniform distribution. At this level, one usually refers to the Vortex Filament Model (VFM), which was originally developed for modeling vortices in superfluid helium Tsubota et al. (2013). It is a quasiclassical model, where one writes Newton’s equations for quantum vortices (presented as filaments in 3D or points in 2D) that interact with all other elements of the system. The drag force exerted on vortices by the impurities is the origin of the mutual force that appears in the hydrodynamical model. Increasing the resolution further, we start to resolve the building blocks of the neutron star matter: neutrons, protons, and electrons (and possibly other elementary particles deep in the star’s core). At this level, one needs to revert to the microscopic description grounded in many-body quantum mechanics. The quantized vortices turn out to be a collective flow of neutrons around some line in space, while impurities correspond to regions where protons (together with neutrons) concentrate into various structures Negele and Vautherin (1973); J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018), usually into nuclei-like objects, but more complex structures are also allowed Caplan and Horowitz (2017). The effective interaction between the flow of neutrons and the bound state of protons is the origin of the drag force that enters the mesoscopic models.
@@
where $Q$ is the selected collective variable. Such a form is achievable once the adiabatic limit can be recovered in the system, which means that the terms describing couplings to other degrees of freedom can be neglected Baranger and Vénéroni (1978); Villars (1977); Goeke et al. (1981). The first term is the kinetic term that is quantified by the mass parameter $M_{Q}$, and $V$ stands for the potential energy. The choice of the collective variable is motivated by physical intuition, although a more precise definition is also possible. In the latter case, one searches for the collective path (or space), which is maximally decoupled from the other degrees of freedom (see Washiyama et al. (2024); Wen and Nakatsukasa (2022); Nakatsukasa et al. (2016) and references therein). This can be important in the case where the choice of collective variables is not obvious, as in the case of exotically deformed nuclei forming the pasta phase. Whatever method is applied to identify $Q$, it implicitly assumes that the adiabatic limit can be achieved, in which the motion of the system is governed by Hamiltonian (1).
In this paper, we present a general-purpose method for extracting mass parameters for nuclear systems by means of the time-dependent density functional theory approach (TDDFT), which is considered as an efficient tool to investigate various problems of nuclear dynamics at low energies at the microscopic level Nakatsukasa et al. (2016). In the field of neutron stars, TDDFT has been applied e.g. in examination of nuclear pasta phase Schuetrumpf et al. (2013), entrainment effects in the inner crust Sekizawa et al. (2022), pairing correlations Stetcu et al. (2011) or development of band theory Kashiwaba and Nakatsukasa (2019). While there are methods for extracting the effective mass parameters, as shown e.g. in Umar and Oberacker (2009); Simenel and Umar (2018), most of them rely on various assumptions, and they are typically designed to be applicable for extracting the mass parameters in a vacuum (see, e.g., the review paper Li and Vretenar (2022)). The presented approach is free of these limitations, and we will demonstrate it for the nuclear impurity immersed in superfluid neutron liquid, i.e., for the conditions relevant to the neutron star crust. The calculations will be delivered for mass parameters associated with translational motion and quadrupole vibrations. The selected collective variables are schematically presented in Fig. 1.
@@ II Method
To make the problem more concrete, let us consider a Wigner crystal of neutron-rich nuclei in the crust of a neutron star. These nuclei can oscillate within the crystal, as illustrated in Fig. 1(a). The collective variable $Q$ can then be associated with the impurity’s position $\bm{R}_{\textrm{c.m.}}$, which, in a microscopic description, corresponds to the center of mass of the protons. The potential energy in this system is primarily due to Coulomb interactions between impurities. Even in this relatively simple setting, determining the effective mass of an impurity immersed in neutron matter remains a challenging task. A variety of approaches have been developed to address this problem Magierski (2004); Magierski and Bulgac (2004); Martin and Urban (2016); N. Chamel, D. Page and S. Reddy (2013); Pęcak et al. (2024). The Wigner crystal is also expected to be threaded by quantum vortices. Work G. Wlazłowski, K. Sekizawa, P. Magierski, A. Bulgac and M. McNeil Forbes (2016) demonstrated that the interaction between an impurity and a vortex can induce a quadrupole deformation of the impurity, Fig. 1(b). This suggests that the quadrupole moment $Q_{20}$ could also serve as a relevant collective variable in the effective model. It is also possible that higher-order nuclear multipole moments play a significant role.
The shape parametrization given by
$R(\theta,\phi)=R_{0}\left(1+\sum_{\lambda=0}^{\infty}\sum_{\mu=-\lambda}^{\lambda}\alpha_{\lambda\mu}Y_{\lambda\mu}(\theta,\phi)\right),$ (2)
is commonly used in effective (collective) models to describe surface deformations of an impurity. Here, $R_{0}$ denotes the radius of the undeformed configuration, and the coefficients $\alpha_{\lambda\mu}$ represent collective degrees of freedom associated with shape oscillations. In this framework, each $\alpha_{\lambda\mu}$ corresponds to a collective mode with its own mass parameter. For example, $\lambda=1$ describes dipole (center-of-mass) motion, while $\lambda=2$ corresponds to quadrupole deformation. While this form suggests an explicit impurity with a sharp boundary, we stress that the aim is to extract the related mass parameters from a microscopic approach, where the nucleus and the surrounding neutron medium are treated self-consistently on equal footing. The parametrization is introduced here solely to define and motivate the collective coordinates used in the analysis, and to provide a connection to the effective models where the concept of a shape is well defined.
@@
It is worth mentioning that there are other methods to extract collective mass parameters that are typically defined in a vacuum Li and Vretenar (2022). One example is the adiabatic approximation to the time-dependent Hartree-Fock-Bogoliubov approach (ATDHFB), commonly used to analyze large amplitude collective motion Ring and Schuck (1980) and widely used in nuclear systems Villars (1977); Baranger and Vénéroni (1978); Goeke et al. (1981). The fundamental idea behind this method is that the collective motion of the whole system is slow compared to the independent movement of each of the nucleons that the system consists of. Based on that, one can obtain collective mass, as it has been done in Baran et al. (2011); Washiyama et al. (2024); Wen and Nakatsukasa (2022). Additionally, this approach has various modifications, among which one should point out the standard method called cranking approximation Ring and Schuck (1980); Girod and Grammaticos (1979). Introduced in Inglis (1956) and Thouless and Valatin (1962), the cranking method has become the standard method commonly used in collective motion examination. All these approaches extract the mass parameter from quasiparticle wave function, corresponding to static solutions of time-independent HFB equations ($i\hbar\partial/\partial t\rightarrow E_{\eta,q}$ in Eq. (9)). Despite the fact that they can also be applied in the analysis of large 3D systems, as it has been shown, e.g., in Umar and Oberacker (2009) and Reinhard et al. (1984), such methods encounter difficulties in cases of more complex problems, like nuclei immersed in the neutron matter. These challenges primarily stem from the significant computational cost associated with conventional methods when applied to systems involving a large number of nucleons, typically far exceeding those encountered in isolated nuclei, which usually contain on the order of a hundred nucleons. Moreover, in such systems, it is not feasible to cleanly separate the quasiparticle wave functions into subsets corresponding to the nucleus and the surrounding medium. Instead, all nucleons must be treated on equal footing, as they are strongly coupled through nuclear interactions. The method proposed in this work is designed to address these limitations. It circumvents the need for explicit linearization or construction of collective subspaces by extracting dynamical information directly from the time evolution of suitably chosen one-body observables, which are functionals of the density. While this still requires solving the full time-dependent equations of motion, this step is no longer a practical bottleneck, thanks to the availability of efficient and well-tested numerical implementations of time-dependent Hartree-Fock Maruhn et al. (2014); Abhishek et al. (2024) and time-dependent Hartree-Fock-Bogoliubov Jin et al. (2021); Pęcak et al. (2024) solvers tailored for large-scale nuclear systems. Therefore, the method is universal and can be applied to the region of the inner crust of a neutron star to extract the effective mass parameters associated with various collective degrees of motion.
@@
The method described above is generic and does not depend on specific functional details. In this study, we use the Brussels-Montreal class of Skyrme-type functionals (BSk in short), which are particularly well suited for studies in the context of neutron stars. These functionals are developed using not only experimental data biased on finite-size nuclei but also many-body results such as pairing gaps Cao et al. (2006) that account for infinite nuclear matter. In this way, the bulk properties are incorporated into the model. In this research, the BSk31 parametrization with microscopically deduced pairing has been applied Goriely et al. (2009, 2016), which incorporates realistic pairing gaps and includes self-energy effects. The Brussels-Montreal-Skyrme functional has a generic form:
$E_{\textrm{BSk}}(t)=\int\mathcal{E}_{\textrm{BSk}}(\bm{r},t)\,d\bm{r}$ (11)
@@
$h_{q\sigma\sigma}=-\bm{\nabla}\frac{\delta E}{\delta\tau_{q}}\bm{\nabla}+\frac{\delta E}{\delta\rho_{q}}-\frac{i}{2}\left\{\frac{\delta E}{\delta\bm{j}_{q}},\bm{\nabla}\right\},\\$ (17)
where $\{,\}$ is the anticommutattor.
@@ III.1 Center of mass oscillations
$Z_{\textrm{c.m.}}(t)=\frac{1}{N_{p}}\int z\ \rho_{p}(\bm{r},t)\ d\bm{r},$ (21)
where $N_{p}=\int\rho_{p}(\bm{r})d\bm{r}$ is the number of protons in the cluster, $z$ is the coordinate of the impurity along $z$ axis, $\rho_{p}(\bm{r})$ is the density of protons. By considering the position of protons only, we avoid problems with defining which neutrons belong to the impurity and which to the background. The associated contribution to the mean-field, entering the single particle Hamiltonian, according to Eq. (10), is
$\begin{split}U_{\textrm{p}}=\frac{\delta E}{\delta\rho_{p}}=U_{\textrm{p},\textrm{BSk}}+k\left(Z_{\textrm{c.m.}}(t)-Z_{0}\right)\frac{z}{N_{p}}.\end{split}$ (22)
@@
In order to examine the system described above, we did systematic studies for average nuclear densities $\bar{\rho}$ ranging from $0.002\textrm{~fm}^{-3}$ to $0.051\textrm{~fm}^{-3}$, where the average density is defined as $\bar{\rho}=(N_{n}+N_{p})/V_{WS}$, ie., number of nucleons (protons and neutrons) per Wigner-Seitz cell volume J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018). We have generated initial states, as shown in Fig. 1, by means of solving the static variant of problem Eq. (9). The problem was solved on a spatial mesh of size $N_{x}\times N_{y}\times N_{z}=32\times 32\times 64$ with lattice spacing $\Delta x=\Delta y=\Delta z=1.25$ fm. The associated cutoff energy is fixed through the value of resolution to $E_{\mathrm{cut}}=\hbar^{2}\pi^{2}/(2m_{n}\Delta x^{2})\approx 130$ MeV. In the computation, we have used W-BSK Toolkit, described extensively in Ref. Pęcak et al. (2024), and available as an open-source package via webpage WBS . Next, we have performed time-dependent simulations with modified mean-field potential Eq. (22) with the initial displacement $Z_{0}=1$fm.
@@
Figure 2: a) Position of the center of mass of protons as a function of time for different values of the $k$ parameter. The considered system has average nuclear density $\bar{\rho}=0.015\textrm{~fm}^{-3}$. b) The extracted frequency of oscillations as a function of the $k$ parameter. The dashed line is the fit of Eq. (8), $\omega^{2}=7\cdot 10^{-6}k$.
We note that the simulations are performed with periodic boundary conditions, whereas the potential defined in Eq. (22) does not strictly satisfy this constraint, i.e., $U_{\textrm{p}}(x,y,z)\neq U_{\textrm{p}}(x,y,z+N_{z}\Delta z)$. However, this does not pose an issue in our case, as $U_{\textrm{p}}$ is coupled solely to the proton density, which is spatially localized and effectively vanishes at the boundaries of the simulation box.
It should be noted that the spin–orbit contribution to the nuclear interaction has been neglected in the present work. While this term is known to play a crucial role in the structure and dynamics of finite nuclei, its impact in the neutron star crust is markedly reduced. Previous studies (e.g., J. M. Pearson, N. Chamel, A. Y. Potekhin, A. F. Fantina, C. Ducoin, A. K. Dutta and S. Goriely (2018)) have shown that the spin–orbit interaction has little influence on the results in this regime. Moreover, as discussed in Pęcak et al. (2024), spatial density fluctuations in the crust are expected to be small, and in the limiting case of uniform matter, the spin–orbit term vanishes entirely. Neglecting this term, therefore, represents a reasonable approximation in the present context and, importantly, reduces the computational cost, enabling simulations of sufficiently large volumes relevant for neutron star crust studies. We emphasize that the proposed method is fully general and can be applied regardless of whether the spin–orbit term is included in the underlying energy density functional.
@@
We applied the fitting procedure to extract the effective parameters of the impurity. Fig. 2 shows the center of mass oscillations for different values of control parameter $k$, for selected nuclear density $\bar{\rho}=0.015\textrm{~fm}^{-3}$. As expected, the extracted frequencies of oscillations follow Eq. (8). One can clearly see that the intercept in this case is $\omega_{Z_{\textrm{c.m.}}}=0$. This feature of the center of mass collective degree of freedom is also expected. Since the Hamiltonian is spatially symmetric, i.e., it is invariant under the operation $\bm{r}\rightarrow\bm{r}+\bm{r}_{0}$, shifting the nucleus does not change the energy of the system. Nevertheless, the translational symmetry is broken: the nuclear cluster can form at any place, which is set randomly. The $\omega_{Z_{\textrm{c.m.}}}=0$ mode is then analogous to a zero-energy Goldstone mode associated with the breaking of translational symmetry. This mode does not represent oscillation but rather a shift with a constant velocity that moves the ground state across all degenerate quantum states that are also solutions of the Hamiltonian.
@@
Figure 3: The effective mass $M_{\textrm{c.m.}}$ (in units of neutron mass $m_{n}$) as a function of mean nuclear density $\bar{\rho}$ extracted for small amplitude $A=1$ fm (orangle triangles) and large amplitude $A=1$ fm (green squares) oscillations. We compare the results of the proposed method (orange triangles) with the effective mass obtained recently using a different dynamical scheme (blue diamonds) Pęcak et al. (2024).
@@
Figure 4: a) The damped oscillations of the center of mass of protons for selected density $\bar{\rho}=0.019\textrm{~fm}^{-3}$ (we show that point in panel c), and b) associated absolute value of velocity as a function of time. The velocity is compared to Landau’s velocity. In panel c), we show the extracted damping coefficient $\gamma_{\textrm{c.m.}}$ as a function of average nuclear density.
@@
where $\Delta_{n}$ and ${k_{\mathrm{F}}}_{n}$ are the bulk values of the pairing field and the Fermi wave-vector of neutrons. Note that this quantity is defined for a uniform system, and it is not obvious how well it works for such a non-uniform system as we study. For increased amplitude, we observed damped oscillations as shown in Fig. 4a). From fitting to the Eq. (19) we can accurately reproduce the behaviour of damped oscillations (orange solid line) and extract the damping parameter $\gamma_{\textrm{c.m.}}$. We investigate the absolute value of the velocity of the impurity in Fig. 4b). One can clearly connect the damping with exceeding the Landau velocity by the impurity. The damped oscillator model starts to underestimate the oscillations’ amplitude once the maximum velocity drops significantly below the Landau velocity. A similar mechanism was observed in recent work Pęcak et al. (2024) where the breaking of Cooper pairs was one of the mechanisms of energy dissipation. The last panel of Fig. 4c) shows the damping parameter $\gamma_{\textrm{c.m.}}$ as a function of average nuclear density in the inner crust. This is a monotonically increasing function of the density. In the limit of low density $\bar{\rho}\rightarrow 0$, it corresponds to a nucleus oscillating in the vacuum. Obviously, in this regime, there are no degrees of freedom where the energy can be dissipated, thus no damping is present. Once the medium gets denser, we see that the damping coefficient increases as well. This effect comes from the investigation of Eq. (23) where the pairing field $\Delta_{n}$ vanishes in high densities Cao et al. (2006), while the wave vector in the denominator $k_{\mathrm{F}}$ grows. Note that in this work, we consider only the $^{1}S_{0}$ pairing channel, and neglect other channels for the pairing, such as higher angular momentum channels $^{3}P_{1}$, which is believed to dominate in higher densities.
The case of oscillations with amplitude $A=10$ fm cannot be considered small, as the amplitude becomes comparable to, or even exceeds, the size of the protonic cluster itself, whose radius ranges from $5$ to $10$ fm depending on the density. Nevertheless, extracting the mass parameter from these data still yields reasonable results, see square markers in Fig. 3. Although the method consistently overestimates $M_{\textrm{c.m.}}$ by approximately 101510-15%, the qualitative behavior is well captured. This discrepancy can be viewed as an estimate of the upper bound on the method’s uncertainty.